# Hyperparameter Optimization in Machine Learning

### v2

@@ 8.3 Neural architecture search
Deep learning has revolutionized the traditional approach of hand-crafting features by enabling the automatic learning of features during training. However, the process of creating neural network architectures still depends significantly on manual design, which involves a substantial amount of trial and error. We can extend the idea of hyperparameter optimization to automatically design neural network architectures, a process known as neural architecture search (NAS).
NAS shares similarities with hyperparameter optimization (HPO), but focuses solely on the hyperparameters that define the architecture. However, NAS faces challenges distinct from most HPO problems, as its search spaces are often significantly higher dimensional, exclusively discrete, and of highly conditional structure (e.g., the hyperparameter controlling the number of units in a given layer is only applicable if that layer exists, as controlled by another hyperparameter).
@@ 8.3.1 Search Space
A key question in NAS is how to define a search space to parameterize a neural network architecture. An expressive search space might lead to entirely new and innovate architecture but might be also prohibitively expensive to explore. On the other hand, a too restrictive space will lower the chance to improve upon manually design architectures. We discuss two popular approaches to define NAS search spaces [white-arxiv23]: Marco Search Spaces and Cell Search Spaces.
Macro search spaces.
We assume a pre-defined architecture topology, such as for example residual neural network or a transformer architecture. Macro search spaces consider design decisions to configure the architecture on a macro level, for example whether to use pre- or post-layer normalization or how many heads to use for all multi-head attention layers.
For example, MNASNet [tan-cvpr19] separates an architecture into blocks and for each block optimizes different hyperparameters, such as the convolution type or kernel size. NASBOT [kandasamy-neurips18] defines the entire network as a DAG where nodes perform a certain operation and edges represent direct dependencies between nodes, or in other words, the output of one node is the input of an adjacent node. For each node, the NAS algorithm has to select the operation type and the number of convolutional filters or fully connected nodes.
The advantage of macro search spaces is that they allow parameterizing each block or layer of the architecture individually. However, while this leads to a very expressive space, it will be quite expensive to search.
Cell search space.
Compared to the Macro Search Space, the Cell Search Space considers a block or cell which is a smaller computational unit that is stacked multiple times on top of each other to form the final architecture. zoph:2018:learning_transferable defined each cell as a directed acyclic graph where for each node they select a certain operation to be performed, for example 1x1 convolution or 3x3 convolution and how to place the edges in the graph. To reduce the overall complexity, zoph2016neural distinguished between two different types of cell, which share the same structure but different strides to reduce the spatial dimension of the input.
Instead of assigning operations to each node, liu2018darts assigns operations to every edge of the graph. Nodes that accumulate the feature vectors from all incoming edges [did not understand]. Their proposed method DARTS then uses gradient descent to select an operation for each edge.
ying-icml19 proposed NASBench-101 which consists of a cell search space with a maximum of 7 nodes and 3 operation choices per node. Furthermore, they limited the total number of edges to 9, which led to a total of 400K different architectures. In an expensive offline step, all architectures were evaluated on CIFAR-10 to form a table. This allows running NAS methods by simply looking up the performance of an architecture instead of training it, accelerating the search process by multiple orders of magnitude.
In follow-up work,  dong-iclr20 proposed NASBench-201 which reduced the number of nodes but allowed for more operations per node. While the total number of architectures in this search space is substantially smaller than in NASBench-101, there are no constraints on the total number of edges, which makes it more accessible for different search algorithms. Furthermore, NASBench-201 evaluated each architecture on CIFAR-10, CIFAR-100 and Imagenet-16.
8.3.2 Search strategies
How do we explore the search space? This is particularly challenging in the case of NAS because of the large number of possible architectures in the search space and the expensive nature of training neural network models.
@@ 8.3.2 Search strategies
In one of the first works on NAS, zoph2016neural used a controller based on a recurrent neural network to suggest new architectures. In each iteration, a set of candidate architectures is sampled and evaluated to collect observations, which are then used to update the controller using reinforcement learning. real2019regularized proposed an evolutionary approach that maintains a population of architectures (see Chapter 6). In each iteration, a random architecture is drawn from the population and mutated along a single dimension. While these approaches led to architectures outperforming state-of-the-art models at that point, the original methods required substantial computational resources since each network was trained from scratch.
To accelerate the search process, we can utilize recent advancements from HPO, such as multi-fidelity optimization. For example, there is a body of work [white-neurips21] that tries to predict learning curves of neural networks to stop the training process early.
@@
A more radically different approach to accelerate NAS is so-called weight-sharing NAS. The idea is to define a single super-network that contains all possible architectures in the finite search space. Each path through the super-network represents a single architecture. In the original work by pham-icml18, this method accelerates the search by more than a factor of 1000.
In a similar vein, liu2018darts proposed DARTS, which adds a continuous architecture parameter for every operation at each edge in the super-network. These architecture parameters are jointly optimized with the weights of the neural network in a bi-level fashion using variations of algorithms we described in Chapter 7. The inner loop fits the weights of the super-network to minimize the training loss, and the outer loop optimizes the architecture parameters to minimize the validation loss. After training the super-network, the final architecture is obtained by selecting the $k$-most important operations per edge. This architecture is then usually re-trained from scratch.
@@
Another approach is so-called two-stage NAS [cai-iclr20, yu-eccv20], which first trains the super-network and then runs gradient-free optimization methods, such as random search or evolutionary algorithms, to select the optimal sub-network without any further training. For example, cai-iclr20 proposed the once-for-all model, which provides a single super-network that allows selecting the optimal architecture with weights for a given hardware, minimizing both latency on the target device and validation error.
@@ 8.4 Meta-learning
We discuss here meta-learning since, from a certain point of view, it shares some interesting similarities to HPO. This section only focuses on this point of view, in the case of supervised learning. For a more general overview of meta learning, see e.g., [hospedales:2021:metalearning_neural_networks] and references therein.
@@ 8.4.1 Definition
In meta-learning (also called learning-to-learn), we have a sequence of tasks $\mathcal{T}^{1},\mathcal{T}^{2},\dots,\mathcal{T}^{N}$, where each task $\mathcal{T}^{j}=(p^{j},\mathcal{L}^{j})$ consists of a distribution $p^{j}(x,y)$, $x\in\mathcal{X}$, $y\in\mathcal{Y}^{j}$ over data points and a loss function $\mathcal{L}^{j}$. Note that all tasks share the same input space but their label spaces may differ. Tasks are sampled from a meta distribution $\mathcal{P}$ on tasks $(\mathcal{T})$. Unlike a traditional supervised algorithm (which, given a dataset, will produce a prediction function $h$), a meta-learning algorithm outputs a learning algorithm that is hoped to learn a good prediction function when presented with a new (unseen) task $\mathcal{T}=(p,\mathcal{L})$ sampled from the same meta distribution $\mathcal{P}$. The meta-training phase consists in computing the learning algorithm $A=ML(\mathcal{D}^{1},\dots,\mathcal{D}^{N})$ where $\mathcal{D}^{j}\sim p^{j}$ and $ML$ is the meta-learning algorithm. The meta-test phase, in turn, consists in computing the prediction function using the new task: $h=A(\mathcal{D})$, with $\mathcal{D}\sim p$.
Meta-learning has its own specificity when compared to other settings dealing with several tasks. Compared to multi-task learning [Caruana94, Baxter2000, ruder:2017:overview_multitask_learning], the set of tasks is not fixed in advance. Compared to transfer learning or domain adaptation [pan2009survey, torrey2010transfer, kouw2018introduction], there are several “source” tasks. Typically, we are interested in the scenario where a very small dataset $\mathcal{D}\leavevmode\nobreak\ \sim p$ will be available for the new task (few-shot learning) or even a single example per class (one-shot learning). The main insight is that when the tasks are related, as captured by the distribution $\mathcal{P}$, meta-learning reduces the sample complexity of learning new tasks drawn by the same distribution [Baxter2000, maurer2016benefit]. Note that while HPO tipycally involves the optimization of a “black-box” function, both the meta-training and the meta-test phases in meta-learning more usually involve the optimization of functions with a know structure, enabling the application of gradient-based techniques, as discussed below.
@@ 8.4.2 Meta-learning of common representations
One effective approach to meta-learning consists of using past tasks to build strong representations, which can be later used for few-shot learning on the new task. We briefly discuss this approach here, as it is the most closely related to the HPO problem [franceschi2018bilevel]. As in [Baxter2000], let us introduce the family of hypothesis spaces $\mathbb{H}^{j}=\{\mathcal{H}^{j}\}$ where each $\mathcal{H}^{j}\in\mathbb{H}^{j}$ is a set of functions $h^{j}_{w,\lambda}:\mathcal{X}\mapsto\mathcal{Y}^{j}$ of the form $h^{j}_{w,\lambda}=\gamma_{w}^{j}\circ\Phi_{\lambda}$, i.e., obtained by composing a shared feature learning function $\Phi_{\lambda}:\mathcal{X}\mapsto\mathcal{F}$, parameterized by $\lambda$ (also called meta-parameter), with a task-specific function $\gamma_{w}^{j}:\mathcal{F}\mapsto\mathcal{Y}_{j}$, parameterized by $w$. The set $\mathcal{F}$ is a suitably chosen feature space, for example $\mathcal{F}=\mathbb{R}^{d}$. This type of function composition is common in several deep learning architectures where $\Phi$ is in charge of representation learning while $\gamma$ could be a classification or regression head. The multi-task network introduced in [Caruana94] and the feature learning neural network described in [Baxter2000] are structured in that way, although all the weights $\lambda$ and $w$ are optimized jointly, unlike the problem formulated below. Note that we have deliberately used the notation $\lambda$ for the parameters shared across the tasks, to highlight the fact that they play a role similar to hyperparameters. In order to construct a “response function,” for every source task $\mathcal{T}^{j}$, $j=1,\dots N$, we split the given dataset $\mathcal{D}^{j}$ into a training set $\mathcal{D}_{\text{tr}}^{j}$ and a validation set $\mathcal{D}_{\text{val}}^{j}$ and define
@@
Figure 8.1: Tasks and datasets for meta-learning on Omniglot [lake:2019:omniglot_challenge_3year]. At the end of the meta-training phase, a small training set for the the new task $\mathcal{T}$ is provided and performance is evaluated on test data $\mathcal{D}_{\text{test}}$. A similar figure was shown in [ravi_optimization_2016] for tasks derived from Imagenet.
@@
where the lower level (or inner) objectives $\mathcal{L}^{j}$ are empirical risk minimization on each training tasks, and the upper (or outer) objective is the validation error on the same tasks. If $\Phi$ and $\gamma$ are realized by deep learning modules, the inner problem would be solved by some iterative gradient-based algorithm, implying that the solution would only be found approximately. Moreover, in general, the above problem is not guaranteed to have a solution. In [franceschi2018bilevel] it is shown that a solution is guaranteed to exist if $\gamma$ is linear (i.e., the classification head consists of a single dense layer) and an algorithm based on hypergradients (using the method described in  Section 7.3) is presented to solve the meta-learning problem. A closely related algorithm is proposed in [finn_model-agnostic_2017] although it is not explicitly formulated as a bilevel program and the emphasis is on fine-tuning to the target task an architecture pretrained on the sequence of previous tasks. A solution based on implicit gradients (see Section 7.2) is offered in [rajeswaran2019meta].
Lastly, while we presented meta-learning as shared feature learning, we note that other tasks’ sharing knowledge mechanisms are possible. Notable examples are biased regularization [denevi2018learning, denevi2019learning, balcan2019provable], in which the meta-parameter $\lambda$ describes a common mean among the tasks, hyper-network approaches, in which $\lambda$ is associated to a function that outputs the parameters of a task-specific models [zhao2020meta, zakerinia2024more], and conditional meta-learning [denevi2020advantage, denevi2022conditional, rusu2018meta, wang2020structured], in which $\lambda$ is a conditioning function of side information specific to the task at hand.

### v3

@@ 8.3 Neural Architecture Search
Deep learning has revolutionized the traditional approach of hand-crafting features by enabling automatic feature learning during training. However, the design of neural network architectures still relies heavily on manual effort, involving substantial trial and error.
Following the problem definition described in Section 2.2, we can extend the idea of hyperparameter optimization (HPO) to automate the design of neural architectures — a process known as neural architecture search (NAS) (see also white-arxiv23, elsken2019neural for an overview). NAS is closely related to HPO but focuses specifically on hyperparameters that define the architecture. At the same time, NAS introduces unique challenges not typically encountered in standard HPO tasks. Its search spaces tend to be higher-dimensional, entirely discrete, and often exhibit highly conditional structures (e.g., the number of units in a layer is only relevant if that layer exists, as determined by another hyperparameter). Furthermore, the high computational cost of training modern deep learning models makes NAS especially expensive and calls for even more efficient optimization techniques.
@@ 8.3.1 Search Space
A key question in NAS is how to define a search space to parameterize a neural network architecture. An expressive search space might lead to entirely new and innovate architecture but might be also prohibitively expensive to explore. On the other hand, a too restrictive space will lower the chance to improve upon manually design architectures. We discuss two popular approaches to define NAS search spaces [white-arxiv23]: Macro Search Spaces and Cell Search Spaces.
Macro search spaces
define the design of neural network architectures at a high level, focusing on configurable choices across the overall structure of the model rather than fine-grained details. Instead of searching over every possible low-level configuration, they assume a fixed architectural topology — such as a residual network or transformer — and explore how the major components within that structure can be customized.
These design decisions might include, for instance, whether to apply layer normalization before or after an operation, how many attention heads to use in each transformer layer, or what type of convolution is used in different stages of a CNN. The goal is to balance architectural flexibility with enough structure to keep the search space manageable. For example, MNASNet [tan-cvpr19] divides the architecture into blocks and tunes high-level choices like convolution type and kernel size per block.
Cell search spaces
focus on smaller computational units — called cells — which are repeatedly stacked to form the full network. Unlike macro search spaces that configure existing architectures, cell-based approaches define a compact building block whose internal structure is optimized and then reused throughout the model allowing to find new topologies.
Typically, each cell is represented as a directed acyclic graph (DAG) where nodes correspond to intermediate feature representations and edges denote operations (e.g., convolutions, pooling). The search algorithm then determines the operations and how they are connected within the cell. This abstraction significantly reduces the size of the search space while retaining enough flexibility to discover high-performing architectures.
For instance, zoph:2018:learning_transferable introduced a cell-based NAS approach where each node in the DAG is assigned a specific operation, such as a 1x1 or 3x3 convolution. To manage complexity and enable spatial downsampling, they distinguished between two types of cells: normal cells, which preserve input dimensions, and reduction cells, which downsample feature maps using larger strides.
8.3.2 Neural Architecture Search as Hyperparameter Optimization
We can phrase NAS as a special type of hyperparameter optimization problem $\min_{\lambda\in\Lambda}f(\lambda)$ (see also Section 2.2), where the inputs $\lambda$ define the architecture and $f(\lambda)$ represents the validation error after training the corresponding network.
@@ 8.3.2 Neural Architecture Search as Hyperparameter Optimization
While early work on automatically discovering neural network architectures dates back some time (for example [kenneth-hyperneat]), bergstra_making_2013 were the first to tackle this problem definition using TPE (see Section 4.2.5) to automatically design architectures for convolutional neural networks. Arguably, zoph2016neural coined the term neural architecture search. In their pioneering work, they proposed a controller based on a recurrent neural network to suggest new architectures for convolutional and recurrent neural networks. In each iteration, a set of candidate architectures is sampled and evaluated to collect observations, which are then used to update the controller using reinforcement learning. While they were able to discover architectures that outperformed the state-of-the-art at the time, their proposed method was prohibitively expensive, requiring up to 800 GPUs concurrently. real2019regularized proposed an evolutionary approach that maintains a population of architectures (see Chapter 6). New candidates are sampled uniformly at random from the population and mutated along a single dimension. Afterward, the population is updated by removing the last entry.
Inspired by the success of Bayesian optimization in HPO, some NAS approaches proposed Bayesian optimization variants to tackle the often high-dimensional and structured search spaces of NAS. kandasamy-neurips18 proposed a Gaussian process variant of Bayesian optimization that measures similarities across architectures using optimal transport approaches. ru-iclr21 argued for combining a Weisfeiler-Lehman graph kernel with a Gaussian process surrogate model instead, which allows for better interpretability of the final architectures. Similar to HPO, these Bayesian optimization approaches usually tend to be faster in terms of convergence speed than evolutionary algorithms such as regularized evolution [ying-icml19, ru-iclr21].
While these approaches led to architectures outperforming state-of-the-art models at the time, the original methods required substantial computational resources since each network was trained from scratch. Nevertheless, these black-box approaches can be quite effective for moderately sized neural networks and smaller search spaces. Furthermore, they can be easily extended to a multi-objective setting, for example, to simultaneously optimize energy consumption and validation performance. Building on recent work on HPO to accelerate the search process, in this setting we can also use more sophisticated approaches, such as multi-fidelity optimization (see Section 5) or transfer learning [feurer2015initializing, wistuba-ieee15, salinas-icml20]. However, due to the ever-growing size of these networks, even these approaches quickly become infeasible.
8.3.3 Weight-Sharing Based Neural Architecture Search
@@
A more radically different approach to accelerate NAS is the so-called weight-sharing NAS. The idea is to define a single super-network that contains all possible architectures in the finite search space. Each path through the super-network represents a single architecture. In the original work by pham-icml18, this method accelerated the search by more than a factor of 1000.
In a similar spirit to gradient-based HPO (see Chapter 7), liu2018darts defined continuous auxiliary parameters for each architectural choice, which are learned by computing the gradient of the validation loss. This resembles the bi-level optimization problem described in Section 2.2; however, here the outer objective optimizes the auxiliary architectural parameters instead of other hyperparameters, and the inner objective optimizes the parameters of the neural network. To reduce computational overhead, liu2018darts proposed DARTS — a gradient-based NAS method — which alternates after each step between the inner and outer objectives. After training the super-network, the best architecture is selected based on the shared weights and then re-trained from scratch. However, several papers [li-uai20, yang-iclr20] reported that this formulation heavily relies on the search space and does not yield better results than just randomly sampling architectures.
@@
Two-stage weight-sharing NAS [cai-iclr20, yu-eccv20] first trains the super-network by updating individual subnetworks in each iteration. After training the super-network, we compute the validation error $f(\lambda)$ by performing a single pass over the validation data. This substantially reduces the computational cost and allows us to apply hyperparameter optimization techniques, such as Bayesian optimization (see Chapter 4) or evolutionary algorithms. For example, cai-iclr20 proposed the once-for-all model, which provides a single super-network that allows selecting the optimal architecture with weights for a given hardware, minimizing both latency on the target device and validation error.
Compared to black-box approaches, weight-sharing methods can be applied to higher-dimensional search spaces and neural networks with a larger parameter count. However, they require more substantial changes to the neural network implementation and therefore cannot be straightforwardly applied to existing codebases. Additionally, they are not easily applicable to more sophisticated cell search spaces defined using a context-free grammar [ericsson-neurips24]. While weight-sharing based methods are popular in the NAS literature, it is not straightforward to apply these techniques to HPO problems, because they only control the capacity of the network and not the training process itself.
@@ 8.4 Meta-Learning
We discuss here meta-learning since, from a certain point of view, it shares some interesting similarities to HPO. This section only focuses on this point of view, in the case of supervised learning. For a more general overview of meta-learning, see for example the survey by hospedales:2021:metalearning_neural_networks and references therein.
@@ 8.4.1 Definition
In meta-learning (also called learning-to-learn), we have a sequence of tasks $\mathcal{T}^{1},\mathcal{T}^{2},\dots,\mathcal{T}^{N}$, where each task $\mathcal{T}^{j}=(p^{j},\mathcal{L}^{j})$ consists of a distribution $p^{j}(x,y)$, $x\in\mathcal{X}$, $y\in\mathcal{Y}^{j}$ over data points and a loss function $\mathcal{L}^{j}$. Note that all tasks share the same input space, but their label spaces may differ. Tasks are sampled from a meta distribution $\mathcal{P}$ over tasks. Unlike a traditional supervised algorithm (which, given a dataset, will produce a prediction function $h$), a meta-learning algorithm outputs a learning algorithm that is hoped to learn a good prediction function when presented with a new (unseen) task $\mathcal{T}=(p,\mathcal{L})$ sampled from the same meta distribution $\mathcal{P}$. The meta-training phase consists in computing the learning algorithm $A=ML(\mathcal{D}^{1},\dots,\mathcal{D}^{N})$ where $\mathcal{D}^{j}\sim p^{j}$ and $ML$ is the meta-learning algorithm. The meta-test phase, in turn, consists in computing the prediction function using the new task: $h=A(\mathcal{D})$, with $\mathcal{D}\sim p$. The concept is illustrated in Figure 8.1.
Meta-learning has its own specificity when compared to other settings dealing with several tasks. Compared to multitask learning [Caruana94, Baxter2000, ruder:2017:overview_multitask_learning], the set of tasks is not fixed in advance. Compared to transfer learning or domain adaptation [pan2009survey, torrey2010transfer, kouw2018introduction], there are several “source” tasks. Typically, we are interested in the scenario where a very small dataset $\mathcal{D}\penalty 10000\ \sim p$ will be available for the new task (few-shot learning) or even a single example per class (one-shot learning). The main insight is that when the tasks are related, as captured by the distribution $\mathcal{P}$, meta-learning reduces the sample complexity of learning new tasks drawn by the same distribution [Baxter2000, maurer2016benefit]. Note that while HPO typically involves the optimization of a “black-box” function, both the meta-training and the meta-test phases in meta-learning more usually involve the optimization of functions with a know structure, enabling the application of gradient-based techniques, as discussed below.
@@ 8.4.2 Meta-learning of Common Representations
One effective approach to meta-learning consists of using past tasks to build strong representations, which can be later used for few-shot learning on the new task. We briefly discuss this approach here, as it is the most closely related to the HPO problem [franceschi2018bilevel]. As in [Baxter2000], let us introduce the family of hypothesis spaces $\mathbb{H}^{j}=\{\mathcal{H}^{j}\}$ where each $\mathcal{H}^{j}\in\mathbb{H}^{j}$ is a set of functions $h^{j}_{w,\lambda}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\mapsto\mathcal{Y}^{j}$ of the form $h^{j}_{w,\lambda}=\gamma_{w}^{j}\circ\Phi_{\lambda}$, i.e., obtained by composing a shared feature learning function $\Phi_{\lambda}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\mapsto\mathcal{F}$, parameterized by $\lambda$ (also called meta-parameter), with a task-specific function $\gamma_{w}^{j}\mathrel{\mathop{\ordinarycolon}}\mathcal{F}\mapsto\mathcal{Y}_{j}$, parameterized by $w$. The set $\mathcal{F}$ is a suitably chosen feature space, for example $\mathcal{F}=\mathds{R}^{d}$. This type of function composition is common in several deep learning architectures where $\Phi$ is in charge of representation learning while $\gamma$ could be a classification or regression head. The multitask network introduced in [Caruana94] and the feature learning neural network described in [Baxter2000] are structured in that way, although all the weights $\lambda$ and $w$ are optimized jointly, unlike the problem formulated below. Note that we have deliberately used the notation $\lambda$ for the parameters shared across the tasks, to highlight the fact that they play a role similar to hyperparameters. In order to construct a “response function,” for every source task $\mathcal{T}^{j}$, $j=1,\dots N$, we split the given dataset $\mathcal{D}^{j}$ into a training set $\mathcal{D}_{\text{tr}}^{j}$ and a validation set $\mathcal{D}_{\text{val}}^{j}$ and define
@@
Figure 8.1: Tasks and datasets for meta-learning on Omniglot [lake:2019:omniglot_challenge_3year]. At the end of the meta-training phase, a small training set for the new task $\mathcal{T}$ is provided and performance is evaluated on test data $\mathcal{D}_{\text{test}}$. A similar figure was shown in [ravi_optimization_2016] for tasks derived from Imagenet.
@@
where the lower level (or inner) objectives $\mathcal{L}^{j}$ are empirical risk minimization on each training tasks, and the upper (or outer) objective is the validation error on the same tasks. If $\Phi$ and $\gamma$ are realized by deep learning modules, the inner problem would be solved by some iterative gradient-based algorithm, implying that the solution would only be found approximately. Moreover, in general, the above problem is not guaranteed to have a solution. In [franceschi2018bilevel] it is shown that a solution is guaranteed to exist if $\gamma$ is linear (i.e., the classification head consists of a single dense layer) and an algorithm based on hypergradients (using the method described in Section 7.3) is presented to solve the meta-learning problem. A closely related algorithm is proposed in [finn_model-agnostic_2017], although it is not explicitly formulated as a bilevel program and the emphasis is on fine-tuning to the target task an architecture pretrained on the sequence of previous tasks. A solution based on implicit gradients (see Section 7.2) is offered in [rajeswaran2019meta].
Lastly, while we presented meta-learning as shared feature learning, we note that other tasks’ sharing knowledge mechanisms are possible. Notable examples are biased regularization [denevi2018learning, denevi2019learning, balcan2019provable], in which the meta-parameter $\lambda$ describes a common mean among the tasks, hyper-network approaches, in which $\lambda$ is associated to a function that outputs the parameters of a task-specific model [zhao2020meta, zakerinia2024more], and conditional meta-learning [denevi2020advantage, denevi2022conditional, rusu2018meta, wang2020structured], in which $\lambda$ is a conditioning function of side information specific to the task at hand.